OpenStudy (anonymous):

Which of the following situations describes a linear relationship? A. The temperature decreases by 10% every day. B. A plant's height triples each week. C. A person walks three feet every second. D. A snowball doubles in size every ten seconds. Which of the following situations describes a linear relationship? A. The temperature decreases by 10% every day. B. A plant's height triples each week. C. A person walks three feet every second. D. A snowball doubles in size every ten seconds.

OpenStudy (schrodinger):

Put it in terms of mathematics: The first one, your variable is T. It decreases by 10%, or is essentially the previous value times .9, every single term. That being said, is the interval it decreases in going to be constant? For the plant, every term is basically the previous value times 3. It will increase almost exponentially. Again, will the interval be constant? The person who is walking is walking at the same speed, with no change in rate. The snowball, again, doubles every ten seconds, but it's increasing. The first time it will be two times the original value, the second "time" it will be quadruple. One of these things is not like the other...

OpenStudy (schrodinger):

(A: 1*.9 = .9... -- .9*.9 = .8, etc.) (B: 1*3 = 3... 3*3 = 9..., etc.) (C: 3*1 = 3... -- 3*2 = 6, etc.) (D: 1*2 = 2... -- 2*2 = 4, etc.)